TIME DOMAIN THz SPECTROSCOPY OF INTRACOMPLEX RYDBERG-RYDBERG TRANSITIONS IN CaF

\underline{VLADIMIR S. PETROVIC}, EMILY E. FENN, and ROBERT W. FIELD, Massachusetts Institute of Technology, Chemistry Department 77 Massachusetts Avenue, Cambridge MA 02139.

In order to develop novel experimental schemes for the study of the Rydberg states in CaF using extraordinarily large dipole moments for intracomplex transitions, a time domain THz spectrometer is being set up. Initial stages of the experiment include rotational time domain THz Spectroscopy of small gas-phase molecules, like CH_{3}F and H_{2}CO, in static cell. In addition, we are employing polarization effects and Stark effect as assignment tools of state-resolved optical-optical double resonance spectra of jet cooled CaF.